# Value-Based Rationales Improve Social Experience: A Multiagent Simulation Study

### v1

@@ 2 Related Work
Whereas other works construct explanations with values [33] or make decisions based on values [8], our focus is on building suitable rationales and investigating how value-aligned rationales shape decision-making and influence social experience, especially privacy. Previous research considers only the causal links between the behaviors and each contextual factor but not the importance of the factors. Is it essential to include every single related factor? Numerous factors may come into play in real-world situations, yet people typically do not provide or need to provide exhaustive explanations to account for every one of these factors.
@@ 3.1 Schematics of an Exanna Agent
the importance of values in one context [23]. We store each value importance $V_{context}$ in a tuple where numbers add up to 1. $v_{i}$ denotes the weight of one value in one value importance ($v_{i}\in V_{context}$) where $0\leq v_{i}\leq 1$ and $\sum_{i=1}^{n}v_{i}=1$. We treat each $\langle V_{context}\rangle$ as an attribute and store the corresponding value importance as its binding. For instance, an agent with value importances $V=\{V_{pandemic}=\{v_{health}=0.6$, $v_{privacy}=0.4\};V_{normal}=\{v_{health}=0.4$, $v_{privacy}=0.6\}\}$ indicates that the agent values health over privacy during a pandemic but the opposite in a normal context.
@@ 3.3 Interaction and Decision Making
Algorithm 1 Decision-making for an Exanna agent
@@
13:        Observe agent $observer$’s action $a_{observer}$
14:        if Receive rationales $Rat_{observer}$ from agent $observer$ then
15:           Update beliefs $b_{t}$ based on $Rat_{observer}$
16:        end if
17:        /* Generate sanctions based on beliefs and rationales */
18:        $sanction_{actor}$ = EvalRationale($Rat_{actor}$ if any, $a_{actor}$, $b_{t}$)
19:     end if
20:     /* Agents learn from reward and sanction */
21:     learn($b_{t}$, $a_{actor}$, $r_{t}+sanction_{actor}$, $b_{t+1}$)
22:  end for
@@ 3.5 Rationale Evaluation
Input: Rationales $Rat$, Observed action $a_{actor}$, Beliefs $b_{t}$
@@
11:  if $act$ contains $a_{actor}$ then
@@ 4.3 Types of Societies
Agents share their decision rules along with selected information that aligns with values of agents present as rationales.
@@ 6 Conclusions, Limitations, and Directions
We thank the anonymous reviewers for their helpful comments. ST and MPS thank the NSF (grant IIS-2116751) for partial support for this research. NA acknowledges partial support from the UKRI EPSRC grant EP/Y028392/1: AI for Collective Intelligence (AI4CI).

### v2

@@ 2 Related Work
Whereas other works construct explanations with values [33] or make decisions based on values [8], our focus is on building suitable rationales and investigating how value-aligned rationales shape decision making and influence social experience, especially privacy. Previous research considers only the causal links between the behaviors and each contextual factor but not the importance of the factors. Is it essential to include every single related factor? Numerous factors may come into play in real-world situations, yet people typically do not provide or need to provide exhaustive explanations to account for every one of these factors.
@@ 3.1 Schematics of an Exanna Agent
The importance of values in one context [23]. We store each value importance $V_{context}$ in a tuple where numbers add up to 1. $v_{i}$ denotes the weight of one value in one value importance ($v_{i}\in V_{context}$) where $0\leq v_{i}\leq 1$ and $\sum_{i=1}^{n}v_{i}=1$. We treat each context as an attribute and store the corresponding $\langle V_{context}\rangle$ as its binding. For instance, an agent with value importances $V=\{V_{pandemic}=\{v_{health}=0.6$, $v_{privacy}=0.4\};V_{normal}=\{v_{health}=0.4$, $v_{privacy}=0.6\}\}$ indicates that the agent values health over privacy during a pandemic but the opposite in a normal context.
@@ 3.3 Interaction and Decision Making
Algorithm 1 Decision making for an Exanna agent
@@
13:        Receive sanction $sanction_{observer}$ from $observer$
14:        Observe agent $observer$’s action $a_{observer}$
15:        if Receive rationales $Rat_{observer}$ from agent $observer$ then
16:           Update beliefs $b_{t}$ based on $Rat_{observer}$
17:        end if
18:        /* Generate sanctions based on beliefs and rationales */
19:        $sanction_{actor}$ = EvalRationale($Rat_{observer}$, $a_{observer}$, $b_{t}$)
20:     end if
21:     /* Agents learn from reward and sanction */
22:     learn($b_{t}$, $a_{actor}$, $r_{t}+sanction_{observer}$, $b_{t+1}$)
23:  end for
@@ 3.5 Rationale Evaluation
Input: Rationales $Rat$, Observed action $a_{observer}$, Beliefs $b_{t}$
@@
11:  if $act$ contains $a_{observer}$ then
@@ 4.3 Types of Societies
Agents share their decision rules along with selected information that aligns with the values of agents present as rationales.
@@ 6 Conclusions, Limitations, and Directions
We thank the anonymous reviewers for their helpful comments. ST and MPS thank the NSF (grant IIS-2116751) for partial support for this research. NA acknowledges support from the UKRI EPSRC grants EP/Y028392/1: AI for Collective Intelligence (AI4CI) and EP/S022937/1: UKRI Centre for Doctoral Training in Interactive AI.